OpenStudy (anonymous):

Factor the polynomial: x^2 + 12x + 11

OpenStudy (anonymous):

x^2 + 11x + x + 11 = x(x+11) +1(x+11) = (x+1)(x+11)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

hi.... sweetvon1 i want to chat with u.......